OpenStudy (anonymous):

Can anyone walk me through this question? Einstein's formula for relativity is E = mc2, where E = energy, m = mass, and c = the speed of light. Solve E = mc2 for c. A. c = square root of quantity E over m B. c = E over m C. c = square root of quantity m over E D.c = m over E

OpenStudy (anonymous):

c= (E/m)^1/2 I guess you can figure now

OpenStudy (anonymous):

That didnt help me at all._.

OpenStudy (perl):

\[\Large{ e=mc^2 \\~\\ \frac{e}{m}=c^2 \\~\\ \sqrt{\frac{e}{m}}=c }\]

OpenStudy (anonymous):

chelsey is offline so idk why your still here im her bro

OpenStudy (nincompoop):

Einstein is saying that mass, the speed of light and energy are all related in that equation.

OpenStudy (anonymous):

she gone can you help me?

OpenStudy (nincompoop):

Theoretical physicist Brian Greene can do a better job :)

OpenStudy (anonymous):

can you help me shes gone

OpenStudy (anonymous):

nevermind

OpenStudy (perl):

Informative video

OpenStudy (perl):

videos*

OpenStudy (astrophysics):

Perl quit hogging the fun questions!

OpenStudy (astrophysics):

=P

OpenStudy (perl):

:)

OpenStudy (anonymous):

Lol your not my bro